OpenStudy (anonymous):

what is b/4+5=14 i cannot do fractions now someone help please asap

OpenStudy (anonymous):

first move 5 to the other side of the equation by subtracting 5 from each side. \[b/4 + 5 - 5 = 14-5\] Then the equation will look like this. \[b/4 = 9\] then you need to get b by itself so multiply both sides by 4. \[b/4 * 4 = 9*4\] I've taken it that far for you, I think you can understand the rest.

OpenStudy (anonymous):

howd u get 5on the other side

OpenStudy (anonymous):

When you are working with equations, what you do to one side of the equals sign you must also do to the other side. So to get rid of 5 on the left you subtract 5 from the left side. But that in turn will also subtract it from the right side of the equation.

OpenStudy (anonymous):

oo ok and uhm do i divide b/4 on both sides?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

no to get rid of division, what is the opposite? Multiplication. so to get rid of something that is divided on one side multiply it. To get rid of something multiplied on one side, divide it. The same works for subtraction and addition.